A 1,200-kg car is traveling on an icy horizontal road at 30 m/s. The maximum rate at which the car can decelerate without sliding on the ice is 1.2 m/s2; reaching this value will immediately result in sliding uncontrollably with a reduced sliding friction of 700 N.

If the driver panics and slams hard on the brakes, then how far will the car slide before stopping? Answer in km.
Answer #1

friction = 700 N

deceleration = F/m = 700 / 1200 = 0.5833 m/s^2

using the equation of linear motion we have

v^2 - u^2 = 2aS

0 - 30^2 = 2*-0.5833*S

S = 771.4 m

so the car slides 771.4 m